# Hölder regularity of Siciak-Zaharjuta extremal functions on compact Hermitian manifolds

### 02 Apr 2026 v1

@@ 5.3. Proof of the third Theorem
Mcshane extension [Mc34] extend a Hölder continuous function from a compact subset to $X$ the same Hölder exponent and the same Hölder constant. Since $X$ is compact, each connected component of $X$ has a finite diameter with respect to the Riemannian distance $d_{\omega}$. Therefore, for a given bounded subset $\{\phi_{\gamma}\}_{\gamma\in\Gamma}$ in $C^{\mu^{\prime}}(K)$, each $\phi_{\gamma}$ can be extended to a function $\tilde{\phi}_{\gamma}$ in $C^{\mu^{\prime}}(X)$ so that $\{\tilde{\phi}_{\gamma}\}_{\gamma\in\Gamma}$ is bounded $C^{\alpha}(X)$. Let $\mu^{\prime\prime}:=(\mu^{\prime})^{2}/(\mu^{\prime}+2+q)$. The direct computation gives

### 05 Apr 2026 v2

@@ 5.3. Proof of the third Theorem
[Mc34], Mcshane extension, extends a Hölder continuous function from a compact subset to $X$ with the same Hölder exponent and the same Hölder constant. Since $X$ is compact, each connected component of $X$ has a finite diameter with respect to the Riemannian distance $d_{\omega}$. Therefore, for a given bounded subset $\{\phi_{\gamma}\}_{\gamma\in\Gamma}$ in $C^{\mu^{\prime}}(K)$, each $\phi_{\gamma}$ can be extended to a function $\tilde{\phi}_{\gamma}$ in $C^{\mu^{\prime}}(X)$ such that $\{\tilde{\phi}_{\gamma}\}_{\gamma\in\Gamma}$ is bounded in $C^{\alpha}(X)$. Let $\mu^{\prime\prime}:=(\mu^{\prime})^{2}/(\mu^{\prime}+2+q)$. The direct computation gives